# Inner-Probe: Discovering Copyright-related Data Generation in LLM Architecture

### v2

@@ Abstract
Index Terms:
@@
I Introduction
Following the release of GPT-3 [1], LLM-driven applications have gained increasing attention in both industry and academia. As transformer architectures mature, the demand for high-quality data has become urgent [2]. However, a major security concern that hinders the establishment of the LLM community is the challenging task of protecting dataset copyright. Specifically, the current model providers lack the capability to identify which datasets contribute most to a response. This inability makes data holders hesitant to expose their datasets to LLM training since copyrighted datasets are of high quality but no credit is given to themselves. Data erasure is the direct method to stop LLM generating copyrighted texts, but the need for finetuning in LLM unlearning approach is computationally heavy and unreliable [3, 4, 5, 6]. Conversely, if model providers could pinpoint the contributions of datasets to a response, a beneficial transactional framework between the data holder and LLM customer could be established [7].
@@ I Introduction
The contribution analysis of sub-datasets to LLM outputs can be framed as a data impact assessment problem, offering a broader application beyond copyrighted data. Typically, mathematical, gradient-based, and prompt-based methods are employed for this task. Mathematical methods adopt a probabilistic approach to compute the similarity between texts and each sub-dataset, such as K-Nearest Access-Free (KNAF) [8, 9]. However, these studies mainly focus on theoretical proof based on strongly ideal assumptions with unknown parameters, which is hard to deploy in practice. Gradient-based machine learning methods, e.g., Shapley values, can be fine-grained into token or data sample level, but it is hard to be directly employed in LLMs because of huge parameters and unrealistic computations [10, 11]. Leveraging the reasoning capability of LLMs, prompt engineering methods aim to ask ‘what the response is based on’, emphasizing factual grounding in cited documents rather than identifying styles or sub-dataset contributions. This approach differs in goals and may amplify hallucinations [12].
For non-copyright content filtering, prior studies have explored watermarking, textual similarity, and prompt-based methods. Data watermarking embeds labels into specific copyrighted data, facilitating source tracing [13, 14, 15]. However, this approach is susceptible to attacks [16]. The token sampling method, which integrates watermarking into the LLM decoding stage, demonstrates greater robustness [17], but the requirement of finetuning LLMs imposes substantial computational overhead [18, 19]. Textual similarity approaches build copyright detectors based on lexical overlap, which uses metrics like LCS and Rouge-L [20], or contextual language models like BERT [21]. However, these methods primarily focus on textual content rather than the text generation process during LLM decoding. This limitation means they may miss cases where copyright-infringing texts depict entirely different scenarios while sharing high-level patterns that textual analysis alone cannot effectively capture. From a prompt-based perspective, previous studies have examined whether specific copyrighted data is included in training datasets [22, 6], using the memorization abilities of LLMs. However, these methods are limited to analyzing source datasets and fail to address copyright issues in LLM-generated outputs.
@@
  • We conduct extensive experiments to validate the effectiveness of Inner-Probe in real-world applications. For the copyrighted sub-dataset contribution identification task, Inner-Probe achieves an accuracy of over 94.9% with a training time of less than 1.5 hours. For the non-copyrighted content filtering task, Inner-Probe attains an accuracy of 93.83% and an AUC score of 0.954 in the Pile dataset [23].
II Background
II-A Limitations of current copyright detection
@@
Before LLMs, techniques for copyright detection focus on learning-based text matching, which relies on small models. Text matching utilizes text similarity methods to match copyrighted texts for those fragments. Statistical one calculates word frequency using TF-IDF and Simhash [24]. However, it ignores text semantics, and is limited to pair-wise comparison, requiring separate computations for each document pair. Deep learning ones use SVM with hand-crafted features [25], autoencoder [26] and bi-directional LSTM (BiLSTM) [27, 28] to capture deep semantic information. Whereas, such methods are only tailored for specific text types and are prone to overfitting, such as newswire stories [26] or comics [29]. More importantly, the definition of copyright infringement is limited to simple text manipulations, including random crop and back-translation of original copyrighted texts [29]. This fails to align with copyright legal definitions, and is inadequate for LLM-generated texts, which has complex variations and adaptations of original copyrighted works.
Limitations of existing methods on LLM copyright detection Currently, most LLM copyright detection frameworks focus on identifying whether a text exists in the entire training database [6, 22]. These methods usually utilize the memorization effects in LLMs, which overemphasize sub-datasets that do not present copyright issues. This results in an incomplete approach, as a comprehensive real-world copyright analysis framework has not yet been fully defined. Additionally, existing studies are constrained by limited context lengths (typically up to 200 tokens) and focus primarily on datasets like books and news [30]. Thus, there is a need for a more concrete LLM copyright framework and a broader testing environment that encompasses diverse content and longer context lengths.
II-B Copyright Protection
@@
Research shows that LLMs can memorize training datasets, leading to efforts to make LLMs forget the data. This method is categorized into parameter-tuning and parameter-agnostic ones. Parameter-tuning involves fine-tuning such as data sharding [31], gradient-based finetuning with reverse loss [32], knowledge distillation [3], additional learnable layer [33], or parameter-efficient module operation [34]. However, these methods can inadvertently affect unrelated information due to knowledge conflicts. To alleviate such fine-tuning costs, the parameter-agnostic method utilizes in-context learning, but the effect is limited to single conversations [35]. Most importantly, using a Min-k probability attack, the unlearning methods prove unreliable [6].
@@
Current studies attempt to prevent LLMs from generating copyrighted content by simple keyword detection and lexical metrics. For example, modifying LLM logits to avoid generating copyright contents by an N-Gram model during decoding [36]. However, these methods remain vulnerable to malicious queries, such as the jailbreak attack that aims to bypass LLM safety checks. White-box attacks use adversarial approaches, using gradient [37], logits [38], or fine-tuning [39], while black-box attacks use template completion [40], rewriting [41] and use LLM to generate attacks [42, 43]. The challenge arises because LLMs function as knowledge databases, making it more likely that copyrighted data is incorporated into generated texts without being directly quoted [44].
II-C Architecture Basic in LLMs
@@ III Copyrights in LLM
III-A LLM-based copyright detection framework
@@ III-B Copyrighted sub-datasets for LLM
III-C Problem formulation for LLM-based copyright detection
@@
III-D Analyzing LLM architecture: Hidden States for Causality
@@
To investigate the architectural information of LLMs beyond textual characteristics, we adopt UMAP, a popular dimensionality reduction method that uses graph layout algorithms to map data into a low-dimensional space [45]. The experiment uses two text sets: (1) 8 classes of original texts from the Pile [23], and (2) corresponding LLM-generated texts through ‘continuation’ prompts by taking original texts as input, simulating a copyright-related generation scenario while maintaining the same class.
@@
Figure 3: Visualization of the statistical differences (UMAP) in hidden states (MHA, FFN) (a) with 8-class input texts from data in the Pile; (b) with input texts that are generated from the pile [23]. For each subfigure in both (a) and (b), Top Row: MHA/FFN output visualization from an LLM (BERT) not trained on the Pile. Bottom Row: MHA/FFN output visualization from an LLM (GPT-series) trained on the Pile across layers.
@@
To learn such a graph from hidden states in LLMs, probabilistic methods are needed. We utilize the Structured Causal Model (SCM) in [46]. We chunk MHA and FFN outputs into matrices, then two causal graphs learn from them. An example result is shown in Fig. 4 . We set the target token (marked in purple) and potential contributing tokens (marked in orange), which are analyzed during learning. We find that MHA constructs comprehensive relationships across tokens, but FFN captures only three token connections, which means MHA shows superior causality analysis for input texts.
@@
IV Inner-Probe Framework
@@ IV-A Overview
IV-B Copyrighted sub-datasets contribution analysis module
In Sec. III-D, we demonstrated that MHA contains significantly more effective information during the text generation process of an LLM model. However, the sequence length of an LLM response is dynamic, making it crucial to design a unified contribution analysis module that can handle varying sequence lengths. In this subsection, we propose two strategies for extracting fixed-length information from MHA with varying lengths, followed by theoretical explanations on how to assess the efficiency of information extraction.
IV-B1 Information extraction and LSTM-based contribution analysis module
@@
In this token extraction method, we directly sample $k$ tokens with interval $n/(k-1)$, where $n$ stands for the total number of tokens in LLM-generated texts, and $Y_{l}$ means MHA outputs in layer $l$. Specifically, when $k=1$, the middle token is sampled. The procedure can be summarized in Eq. 8.
@@
We define pivotal tokens as those with higher representation variance, which shows significance in lightweight LLM inference [47]. As shown in Eq. 9, for layer $l$, we calculate variance $Var(Y_{l}[i])$ for each token position $i$. The ones with top-k highest variance are selected, constructing extracted representation $O_{l}$. As shown in Eq. 9, this method is annotated as Variance (Var).
@@
[…]
@@
As depicted in Fig. 6(a), the classifier captures the layer-wise relationship between extracted MHA outputs, inspired by Observation 1 in section III-D. Therefore, the number of time steps in LSTM matches the layer number in LLM architecture. For each time step, the MLP fusion module first compresses $O_{l}$ into a compact representation, which is then fed into the corresponding LSTM cells. The final output of the LSTM is passed to a two-layer classifier, producing a softmax distribution that represents the contribution score of each copyrighted sub-dataset.
IV-B2 Theoretical Explanations for Different Information Extraction Strategies
@@
IV-C Contrastive-learning based Non-copyright Filtering Module
@@
We achieve non-copyright response filtering by training a module that directly detects non-copyright contexts from $D$. Copyright datasets $D_{c}$ are batched into $\{T_{b}\}_{b=1}^{\mid B\mid}$, and augmented into $\{\tilde{T_{b}}\}_{b=1}^{\mid B\mid}$ using Random Span Masking (RSM) for data augmentation [48, 21]. Thus, the final embeddings of such pairs are denoted as $e_{b}$ and $\tilde{e}_{b}$. These augmented pairs are passed through a frozen LLM to generate layer-wise MHA outputs, which are processed by the LSTM-based feature extractor and the copyrighted sub-datasets contribution analysis module. In this module, the global projector (lightweight MLP) processes the first-layer MLP outputs from the data pairs, and produces two embeddings, $e_{b}$, $\tilde{e}_{b}$. The module is trained using contrastive loss, which encourages similar embeddings of positive pairs while separating embeddings of different samples. Non-copyright issues can then be detected by thresholding in the learned embedding space, as shown in Fig. 6(c). The loss for each query text index $i$ is computed as in Eq. 15, where $\tau$ is the temperature parameter.
@@
We leverage the embedding space learned from copyright data to filter out non-copyright responses based on distance-based measurement. Specifically, we use Mahalanobis distance, a classic distance-based scoring function $d_{M}$ [49], which calculates the distance between a vector point $\mathbf{x}$ and a distribution $\mathbf{R_{Q}}\in\mathbb{R}^{N}$, as defined in Eq. 16. Here, $\mathbf{x}$ represents embedding of LLM response sample, while $\boldsymbol{\mu}$ and $\mathbf{\Sigma}$ represents mean and positive semi-definite covariance matrix of embeddings from distribution of copyrighted responses $\mathbf{R_{Q}}$, and $M$ is the dimensionality of $\boldsymbol{\mu}$. As in Eq. 6, the anomaly threshold $\delta$ determines whether a text is non-copyright. This $\delta$ is set to ensure a 95% true positive rate (TPR) on the training samples.
@@
IV-D Controllable Text Generation Design
A remaining challenge in evaluating the performance of Inner-Probe is collecting the desired LLM responses that suffer copyright issues. Moreover, the ground truth of which specific copyright is infringed in response is also unknown. To address the problem, we propose controllable text generation designs based on copyright-related prompts. Our method first draws inspiration from established copyright infringement criteria for books [50], and extends to analyze broader datasets in the Pile via three generation scenes and text mixture task.

### v3

@@ Abstract
{IEEEImpStatement}
Large Language Models (LLMs) show remarkable text generation capabilities, but their reliance on high-quality copyright datasets has raised significant concerns about infringement. Existing methods include text matching, memorization-based detection, and prompt engineering. However, they fall short in capturing sufficient semantic information with extended contexts, and incur high computational costs. To overcome these limitations, we introduce a lightweight framework that leverages multi-head attention outputs for precise sub-dataset contribution analysis and effective non-copyrighted content filtering. Our approach achieves over 94.9% accuracy in identifying the contributing copyrighted sub-datasets and 93.83% accuracy with an AUC of 0.954 in filtering, substantially outperforming prior methods. This not only mitigates the shortcomings of prior research but also provides a practical, efficient solution for ensuring responsible real-world LLM deployment while protecting intellectual property.
{IEEEkeywords}
@@
1 Introduction
Following the release of GPT-3 [gpt-3], LLM-driven applications have gained increasing attention in both industry and academia. As transformer architectures mature, the demand for high-quality data has become urgent [prove_data_quality]. However, a major security concern that hinders the establishment of the LLM community is the challenging task of protecting dataset copyright. Specifically, the current model providers lack the capability to identify which datasets contribute most to a response. This inability makes data holders hesitant to expose their datasets to LLM training since copyrighted datasets are of high quality but no credit is given to themselves. Data erasure is the direct method to stop LLM generating copyrighted texts, but the need for finetuning in LLM unlearning approach is computationally heavy and unreliable [harry, arxivunlearn, unlearningdowngrade, detecting]. Conversely, if model providers could pinpoint the contributions of datasets to a response, a beneficial transactional framework between the data holder and LLM customer could be established [Cui2024RethinkingUO].
@@ 1 Introduction
The contribution analysis of sub-datasets to LLM outputs can be framed as a data impact assessment problem, offering a broader application beyond copyrighted data. Typically, mathematical, explainability-based attribution, and data traceability methods are developed for this task. Mathematical methods adopt a probabilistic approach to compute the similarity between texts and each sub-dataset, such as K-Nearest Access-Free (KNAF) [follow, 2_privacy]. However, these studies mainly focus on theoretical proof based on strongly ideal assumptions with unknown parameters, which is hard to deploy in practice. Explainability-based attribution methods are separated into fine-tuning-based and prompt-based paradigms [mezzi2025owns]. The former utilizes Shapley values (SHAP) and Local Interpretable Model-Agnostic Explanations (LIME), which explain output behaviors at the token or data sample level [SHAPcite, LIMEcite]. However, it is challenging to be directly employed in LLMs due to the huge parameters and unrealistic gradient computations [Anthropic, sharply]. Leveraging the reasoning capability of LLMs, the prompt-based paradigm aims to ask ‘what the response is based on’, emphasizing factual grounding in cited documents rather than identifying styles or sub-dataset contributions. This approach differs in goals and may amplify hallucinations [hal]. Data traceability refers to tracking the origin, process, and destination of data. Despite studies in large-scale commercial database management [karkovskova2021design, tang2019sac], currently, there is no investigation into the LLM generation process.
For non-copyright content filtering, prior studies have explored generated content identification, textual similarity, and LLM memorization [mezzi2025owns]. Generated content identification includes watermarking and fingerprinting. Data watermarking embeds labels into specific copyrighted data, facilitating insertion-based source tracing [distribution, linguistic, entropy]. However, this approach is vulnerable to attacks, and the requirement of finetuning LLMs imposes huge computational overhead [panaitescu2024can, training1, wasa]. Fingerprinting passively extracts inherent characteristics from content to distinguish it from other similar objects, enabling identification of LLM-generated text. But it mostly refers to model-level copyrights, not copyrighted datasets [llmmap]. Textual similarity approaches build copyright detectors based on lexical overlap, using metrics like LCS and Rouge-L [liu-etal-2024-shield], or contextual language models like BERT [cho2022enhancing]. However, these methods primarily focus on textual content rather than the text generation process during LLM decoding. This limitation means they may miss cases where copyright-infringing texts depict entirely different scenarios while sharing high-level patterns that textual analysis alone cannot effectively capture. Lastly, previous studies have examined whether specific copyrighted data is included in training datasets [decop, detecting], using the memorization abilities of LLMs. However, these methods are limited to analyzing source datasets and fail to address copyright issues in LLM-generated outputs.
@@
  • We conduct extensive experiments to validate the effectiveness of Inner-Probe in real-world applications. For the copyrighted sub-dataset contribution identification task, Inner-Probe achieves an accuracy of over 94.9% with a training time of less than 1.5 hours. For the non-copyrighted content filtering task, Inner-Probe attains an accuracy of 93.83% and an AUC score of 0.954 in the Pile dataset [gao2020pile].
2 Background
2.1 Positioning and taxonomy
At present, there exist four main attribution method classes: (a) content identification (watermarking, fingerprinting), (b) security-based techniques (blockchains, digital signatures, zero-knowledge proofs, synthetic media forensics), (c) explainability methods (fine-tuning and prompt-based), and (d) data traceability (data lineage) [mezzi2025owns].
Accordingly, we position Inner-Probe within data traceability [mezzi2025owns]. The first two classes answer whether specific copyrighted content is present. But they do not quantify how underlying training datasets contribute to a given generation. The latter two classes apply to LLM training data attribution. However, explainability methods often suffer from hallucination or require LLM fine-tuning costs to yield stable results. Meanwhile, existing data-traceability works concentrate on enterprise data management rather than tracing contributions through the LLM generation process [karkovskova2021design, tang2019sac]. To contextualize our evaluation, we also draw on a survey of prompt-based LLM attribution. Following its taxonomy, we include a representative post-generation attribution (explainability) method as one of our baselines [attribution_survey].
2.2 Limitations of current copyright detection
@@
Before LLMs, techniques for copyright detection focus on learning-based text matching, which relies on small models. Text matching utilizes text similarity methods to match copyrighted texts for those fragments. Statistical one calculates word frequency using TF-IDF and Simhash [jin2017research]. However, it ignores text semantics, and is limited to pair-wise comparison, requiring separate computations for each document pair. Deep learning ones use SVM with hand-crafted features [prieto2018supervised], autoencoder [briciu2021autoat] and bi-directional LSTM (BiLSTM) [tavan2021bert, petropoulos2024roberta] to capture deep semantic information. Whereas, such methods are only tailored for specific text types and are prone to overfitting, such as newswire stories [briciu2021autoat] or comics [xin2024text]. More importantly, the definition of copyright infringement is limited to simple text manipulations, including random crop and back-translation of original copyrighted texts [xin2024text]. This fails to align with copyright legal definitions, and is inadequate for LLM-generated texts, which has complex variations and adaptations of original copyrighted works.
Limitations of existing methods on LLM copyright detection Currently, most LLM copyright detection frameworks focus on identifying whether a text exists in the entire training database [detecting, decop]. These methods typically utilize the memorization capabilities of LLMs, which can overemphasize sub-datasets that do not raise copyright issues. This results in an incomplete approach, as a comprehensive real-world copyright analysis framework has not yet been fully defined. Additionally, existing studies are constrained by limited context lengths (typically up to 200 tokens) and focus primarily on datasets like books and news [copyrighttakedown]. Thus, there is a need for a more concrete LLM copyright framework and a broader testing environment that encompasses diverse content and longer context lengths.
2.3 Copyright Protection
@@
Research shows that LLMs can memorize training datasets, leading to efforts to make LLMs forget the data. This method is categorized into parameter-tuning and parameter-agnostic ones. Parameter-tuning involves fine-tuning such as data sharding [liu2024forgetting], gradient-based finetuning with reverse loss [yu2023unlearning], knowledge distillation [harry], additional learnable layer [chen-yang-2023-unlearn], or parameter-efficient module operation [hu2024separate]. However, these methods can inadvertently affect unrelated information due to knowledge conflicts. To alleviate such fine-tuning costs, the parameter-agnostic method utilizes in-context learning, but the effect is limited to single conversations [DBLP:conf/icml/PawelczykNL24]. Most importantly, using a Min-k probability attack, the unlearning methods prove unreliable [detecting].
@@
Current studies attempt to prevent LLMs from generating copyrighted content by simple keyword detection and lexical metrics. For example, modifying LLM logits to avoid generating copyright contents by an N-Gram model during decoding [ippolito-etal-2023-preventing]. However, these methods remain vulnerable to malicious queries, such as the jailbreak attack that aims to bypass LLM safety checks. White-box attacks use adversarial approaches, using gradient [zou2023universal], logits [zhou2024don], or fine-tuning [iclr24oral], while black-box attacks use template completion [deng2024pandora], rewriting [YuanJW0H0T24] and use LLM to generate attacks [casper2023explore, DBLP:conf/camlis/BirchHTSG23]. The challenge arises because LLMs function as knowledge databases, making it more likely that copyrighted data is incorporated into generated texts without being directly quoted [cyphert2023generative].
2.4 Architecture Basic in LLMs
@@ 3 Copyrights in LLM
3.1 LLM-based copyright detection framework
@@ 3.2 Copyrighted sub-datasets for LLM
3.3 Problem formulation for LLM-based copyright detection
@@
3.4 Analyzing LLM architecture: Hidden States for Causality
@@
To investigate the architectural information of LLMs beyond textual characteristics, we adopt UMAP, a popular dimensionality reduction method that uses graph layout algorithms to map data into a low-dimensional space [mcinnes2018umap]. The experiment uses two text sets: (1) 8 classes of original texts from the Pile [gao2020pile], and (2) corresponding LLM-generated texts through ‘continuation’ prompts by taking original texts as input, simulating a copyright-related generation scenario while maintaining the same class.
@@
Figure 3: Visualization of the statistical differences (UMAP) in hidden states (MHA, FFN) (a) with 8-class input texts from data in the Pile; (b) with input texts that are generated from the pile [gao2020pile]. For each subfigure in both (a) and (b), Top Row: MHA/FFN output visualization from an LLM (BERT) not trained on the Pile. Bottom Row: MHA/FFN output visualization from an LLM (GPT-series) trained on the Pile across layers.
@@
To learn such a graph from hidden states in LLMs, probabilistic methods are needed. We utilize the Structured Causal Model (SCM) in [neurips23causal]. We chunk MHA and FFN outputs into matrices, then two causal graphs learn from them. An example result is shown in Fig. 4 . We set the target token (marked in purple) and potential contributing tokens (marked in orange), which are analyzed during learning. We find that MHA constructs comprehensive relationships across tokens, but FFN captures only three token connections, which means MHA shows superior causality analysis for input texts.
@@
4 Inner-Probe Framework
@@ 4.1 Overview
4.2 Copyrighted sub-datasets contribution analysis module
In Sec. 3.4, we demonstrated that MHA contains significantly more effective information during the text generation process of an LLM model. However, the sequence length of an LLM response is dynamic, making it crucial to design a unified contribution analysis module that can handle varying sequence lengths. In this subsection, we propose two strategies for extracting fixed-length information from MHA with varying lengths, followed by theoretical explanations on how to assess the efficiency of information extraction.
4.2.1 Information extraction and LSTM-based contribution analysis module
@@
In this token extraction method, we directly sample $k$ tokens with interval n/(k1)n/(k-1), where $n$ stands for the total number of tokens in LLM-generated texts, and $Y_{l}$ means MHA outputs in layer $l$. Specifically, when $k=1$, the middle token is sampled. The procedure can be summarized in Eq. 8.
@@
We define pivotal tokens as those with higher representation variance, which shows significance in lightweight LLM inference [DBLP:conf/aaai/LeeJKKP24]. As shown in Eq. 9, for layer $l$, we calculate variance $Var(Y_{l}[i])$ for each token position $i$. The ones with top-k highest variance are selected, constructing extracted representation $O_{l}$. As shown in Eq. 9, this method is annotated as Variance (Var).
@@
0:  Strategy $s\in\{\textsc{INTER},\textsc{VAR},\textsc{A-VAR}\}$, Number of selected tokens $k$ per layer
@@
As depicted in Fig. 6(a), the classifier captures the layer-wise relationship between extracted MHA outputs, inspired by Observation 1 in section 3.4. Therefore, the number of time steps in LSTM matches the layer number in LLM architecture. For each time step, the MLP fusion module first compresses $O_{l}$ into a compact representation, which is then fed into the corresponding LSTM cells. The final output of the LSTM is passed to a two-layer classifier, producing a softmax distribution that represents the contribution score of each copyrighted sub-dataset.
4.2.2 Theoretical Explanations for Different Information Extraction Strategies
@@
4.3 Contrastive-learning based Non-copyright Filtering Module
@@
We achieve non-copyright response filtering by training a module that directly detects non-copyright contexts from $D$. Copyright datasets $D_{c}$ are batched into $\{T_{b}\}_{b=1}^{\mid B\mid}$, and augmented into $\{\tilde{T_{b}}\}_{b=1}^{\mid B\mid}$ using Random Span Masking (RSM) for data augmentation [liu2021fast, cho2022enhancing]. Thus, the final embeddings of such pairs are denoted as $e_{b}$ and $\tilde{e}_{b}$. These augmented pairs are passed through a frozen LLM to generate layer-wise MHA outputs, which are processed by the LSTM-based feature extractor and the copyrighted sub-datasets contribution analysis module. In this module, the global projector (lightweight MLP) processes the first-layer MLP outputs from the data pairs, and produces two embeddings, $e_{b}$, $\tilde{e}_{b}$. The module is trained using contrastive loss, which encourages similar embeddings of positive pairs while separating embeddings of different samples. Non-copyright issues can then be detected by thresholding in the learned embedding space, as shown in Fig. 6(c). The loss for each query text index $i$ is computed as in Eq. 15, where $\tau$ is the temperature parameter.
@@
We leverage the embedding space learned from copyright data to filter out non-copyright responses based on distance-based measurement. Specifically, we use Mahalanobis distance, a classic distance-based scoring function $d_{M}$ [lee2018simple], which calculates the distance between a vector point $\mathbf{x}$ and a distribution $\mathbf{R_{Q}}\in\mathbb{R}^{N}$, as defined in Eq. 16. Here, $\mathbf{x}$ represents embedding of LLM response sample, while $\boldsymbol{\mu}$ and $\mathbf{\Sigma}$ represents mean and positive semi-definite covariance matrix of embeddings from distribution of copyrighted responses $\mathbf{R_{Q}}$, and $M$ is the dimensionality of $\boldsymbol{\mu}$. As in Eq. 6, the anomaly threshold $\delta$ determines whether a text is non-copyright. This $\delta$ is set to ensure a 95% true positive rate (TPR) on the training samples.
@@
4.4 Controllable Text Generation Design
A remaining challenge in evaluating the performance of Inner-Probe is collecting the desired LLM responses that suffer copyright issues. Moreover, the ground truth of which specific copyright is infringed in response is also unknown. To address the problem, we propose controllable text generation designs based on copyright-related prompts. Our method first draws inspiration from established copyright infringement criteria for books [wikipediaSubstantialSimilarity], and extends to analyze broader datasets in the Pile via three generation scenes and text mixture task.